OpenStudy (anonymous):

the sum of Hank's age and Albert's age is 60. if Hank's age 8 years from now will be three times Albert's age 4 years ago, how old is Hank now?

OpenStudy (anonymous):

Hank's age --> x Albert's age --> y

OpenStudy (anonymous):

can you write the question in algebra now?

OpenStudy (anonymous):

@LuizaMiller

OpenStudy (anonymous):

Hank=h Albert=a 60-8=52 in 4 years ago a+h=52 3a+8=h OK ?! @LuizaMiller

OpenStudy (anonymous):

X+y=60 3(y-4)=x+8 Hank is 40

OpenStudy (anonymous):

Is it ok PFEH.1999 ???

OpenStudy (anonymous):

E.ali oh thank you very much

OpenStudy (anonymous):

Your welcome !:)

OpenStudy (anonymous):

no no,you did it wrong!

OpenStudy (anonymous):

x + y = 60 , yes it's true!

OpenStudy (anonymous):

I got confused is it right or is it wrong????

OpenStudy (anonymous):

what did you find?hank's age = ?

OpenStudy (anonymous):

40

OpenStudy (anonymous):

yes it's true , sorry... Hank's age is 40 and Albert's age is 20 sorry again :)

OpenStudy (anonymous):

:) it's ok darlin'